OpenStudy (anonymous):

What is another way to write h≠0? my precal teacher doesnt let us do that.

OpenStudy (anonymous):

what exactly are you trying to say?

myininaya (myininaya):

h->0 is another way

OpenStudy (anonymous):

they don't really mean the same thing at all do they?

myininaya (myininaya):

h->0 means we are approaching h as never actually land on h

OpenStudy (anonymous):

\[h\neq 0\] means h is any non-zero number

OpenStudy (anonymous):

that is not the same as taking a limit for sure

myininaya (myininaya):

right i just figure he was talking about limits

OpenStudy (anonymous):

i was solving a problem dealing with difference quotient

OpenStudy (radar):

What does your teacher specify when h does not = zero? I thought that they symbol\[\neq\]

OpenStudy (radar):

was pretty much universal.

OpenStudy (anonymous):

should be x not h

OpenStudy (anonymous):

\[\frac{x^2-4}{x-2}=x+2\text{ if } x\neq 0\] \[\lim_{x\rightarrow 2}\frac{x^2-4}{x-2}=2+2=4\]

myininaya (myininaya):

it x does not equal 2?

myininaya (myininaya):

lol

myininaya (myininaya):

\[\frac{x^2-4}{x-2}=x+2\text{ if } x\neq 2\]

OpenStudy (anonymous):

maybe you should ask your precalc teacher what they want. i have seen the domain of \[\frac{x^2-4}{x-2}\] written as \[(-\infty,2)\cup(2,\infty)\]

myininaya (myininaya):

that is the domain

OpenStudy (anonymous):

which is correct but a waste of a perfectly good \[x\neq 0\]